# Diversity in Fermi/GBM Gamma Ray Bursts: New insights from Machine Learning

### v1

@@ Abstract
Classification of gamma-ray bursts (GRBs) has been a long-standing puzzle in high-energy astrophysics. Recent observations challenge the traditional short vs. long viewpoint, where long GRBs are thought to originate from the collapse of massive stars and short GRBs from compact binary mergers. Machine learning (ML) algorithms have been instrumental in addressing this problem, revealing five distinct GRB groups within the Swift/BAT light curve data, two of which are associated with kilonovae (KNe). We corroborate these five classes by extending this analysis to the Fermi/GBM data using unsupervised ML techniques. These five clusters are well separated in fluence-duration plane, hinting at a potential link between fluence, duration and complexities (or structures) in the light curves of GRBs. Further, we confirm two distinct classes of KN-associated GRBs. The presence of GRB 170817A in one of the two KNe-associated clusters lends evidence to the hypothesis that this class of GRBs could potentially be produced by binary neutron star (BNS) mergers. The second KN-associated GRB cluster could potentially originate from NS-BH mergers. Future multimessenger observations of compact binaries in gravitational waves (GWs) and electromagnetic waves can be paramount in understanding these clusters better.
@@ 1 Introduction
Recent advances in machine learning (ML) and artificial intelligence (AI) are fostering a more detailed understanding of GRBs (Chattopadhyay & Maitra, 2017; Acuner & Ryde, 2018; Jespersen et al., 2020a; Salmon et al., 2022; Steinhardt et al., 2023; Garcia-Cifuentes et al., 2023; Mehta & Iyyani, 2024; Negro et al., 2024). When applied to the analysis of prompt emission light curves, these techniques have revealed hidden patterns within the GRB population, suggesting the existence of subclasses beyond the traditional short–long GRB classification (Jespersen et al., 2020b; Steinhardt et al., 2023; Garcia-Cifuentes et al., 2023; Dimple et al., 2023a).
@@ 2.1 Data Reduction Pipeline
After standardization, we could recover 3516 GRBs from the Fermi/GBM catalog. For the ML algorithm, we used the data in three energy bands, two energy bands of NaI (8–50 keVand 50–300 keV) and one energy band of BGO (300–100 keV) as NaI is more sensitive in the lower energy band, and BGO is more sensitive in the high energy bands. For both datasets, we used binning of 16ms.
@@ 2.2 Clustering using unsupervised learning
We generated the connectivity and density maps to further analyze the resulting low-dimensional representation. Connectivity maps visualize the probability of a connection between two data points, essentially depicting the aforementioned network in a two-dimensional format. On the other hand, density maps, along with the 2-dimensional embeddings, offer additional insights into the data structure. These maps help to understand the distribution and concentration of data points in a low-dimensional space. Finally, we identified clusters within the two-dimensional embedding using the AutoGMM module (Athey et al., 2019), which applies a Gaussian Mixture Model (GMM) to automatically determine the optimal number of clusters present in the data.
Figure 2: GRB locations on fluence–$T_{\rm 90}$ plane: GRBs are color coded based on distinct clusters identified by AutoGMM. Notably, these clusters exhibit clear separation in the fluence-duration plane, implying a potential correlation between light curve features and GRB fluence and duration.
@@ 3.1 Distinct groups within GRBs
The consistency of five distinct clusters across multiple datasets and methodologies suggests that the GRB population exhibits diversity beyond just two classes. To further investigate these clusters, we examined their prompt emission properties. We observed that these clusters exhibit minimal differences in their spectral properties, such as $E_{\rm p}$ and $\alpha$ (See Figure A.1 in Appendix A). To explore any correlation between fluence and duration for these clusters, we plotted the GRBs from the five clusters on the fluence–$T_{\rm 90}$ plane (Figure 2). Interestingly, despite using fluence-normalized light curves for classification, they are well separated in this plane. This suggests a strong correlation between the brightness and complexity of the GRB light curves, a prediction that aligns with the findings of Stern et al. (1997). This also allows the identification of various classes: short-duration bursts with low fluence (cluster 0), intermediate-duration bursts with low fluence (clusters 1 and 3), long-duration bursts with low fluence (cluster 4), and intermediate to long-duration bursts with high fluence (cluster 2). Though the correlation between fluence and duration is well known (Horváth et al., 2001), our results hint at a correlation between light curve shapes and brightness, which may have a deeper connection with the physics of the central engines of GRBs. A detailed study of the implications of this correlation for the central engines of GRBs may be the subject of future work.
@@
Given that our clustering is based solely on the prompt emission light curves, these distinct classes may represent different progenitors, central engines, emission mechanisms, or a combination of these. To date, evidence from optical bumps (SNe/KNe) suggests that the collapse of massive stars (collapsars) or the merging of compact objects (NS/BH/WD) are two possible scenarios for the origin of GRBs. We locate all the candidates for KNe and SNe on the connectivity maps. SN-associated GRBs are spread across the map, as evident from Figure 3. Interestingly, some of these candidates are adjacent to KN-associated GRBs. Li et al. (2023) also reported similarity in the temporal and spectral properties of some of the KN- and SN-associated GRBs. This indicates that despite differences in the progenitors, they might have similar central engines and/or emission mechanisms, resulting in similar light curves. It is possible that collapsars and compact binary mergers may result in the formation of a central engine, such as a hyperaccreting black hole or a magnetar, which could be responsible for launching a relativistic jet (Duncan & Thompson, 1992; Usov, 1992; Popham et al., 1999; Narayan et al., 2001; Metzger, 2019).
@@ 3.2 Two distinct populations of KN-associated GRBs
The results are presented in Figure 3, which interestingly confirms the existence of two clusters of KNe-associated GRBs based on Swift/BAT. Unlike Swift/BAT, Fermi/GBM has observed GRB 170817A, known to be a BNS merger from GW observations (Abbott et al., 2017) with a confirmed detection of a KN. Therefore, we find evidence favouring the conjecture of Dimple et al. (2023a) that the low fluence, small duration cluster may correspond to BNS mergers. Similarly, GRBs 170817A and 160821B are closely located in the same cluster, implying they could possibly originate from the same kind of progenitor systems. Dimple et al. (2023a) argued that the second KNe-associated cluster could potentially be arising from NS-BH mergers. As GRB 230307A is not detected by Swift/BAT, the proximity of GRBs 211211A and 230307A in the Fermi/GBM map suggests the similarity in their light curves and the progenitors. They are lying in a close vicinity in the low-dimensional embeddings obtained using deep learning on Fermi/GBM data (Negro et al., 2024). An independent study of the temporal and spectral properties of these GRBs also suggests that they might have the same origin (Peng et al., 2024). However, multimessenger observation of a NS-BH system is necessary to confirm the prediction that GRBs in this cluster are products of NS-BH mergers. For the typical NS masses, the BH must be of mass $\sim 3M_{\odot}$ for the NS-BH to produce a GRB. Recent observation of an NS-BH merger GW230529 (Abac et al., 2024) with a BH between $2.5-4.5M_{\odot}$ indicates there may be a population of this low mass BH in the universe. With all these, the first multimessenger NS-BH merger will revolutionize our understanding of compact objects and GRBs.
@@
    Of these five clusters, two distinct clusters contain KNe-associated GRBs. One cluster contains GRB 170817A, which is associated with a BNS merger. This strengthens the case for these two clusters to be associated with BNS and NS-BH mergers as proposed by Dimple et al. (2023a). However, there are a couple of SNe-associated GRBs in these clusters which may possibly indicate the universality of the central engines of GRBs regardless of their progenitors.
@@
    Despite the use of fluence-normalized light curves, the ML-based clusters are well-separated in the fluence-$T_{\rm 90}$ plane, hinting that fluence played a pivotal role in clustering and may have a strong correlation with the duration and complexities or structures in the light curve of the GRBs.
We next discuss the implications of these findings. Even though the Fermi and Swift datasets contain mostly distinct events, with only a small overlap, their clustering patterns are remarkably similar. This consistency highlights the robustness of the ML-based classification. Further, KNe-associated GRBs are tightly clustered at two locations in both datasets. These clustering patterns can be used for early identification of such GRBs. This may be achieved by reanalyzing the existing data with the light curve of a new GRB in question and asking whether it lies in one of the two clusters. Such an algorithm can be developed and applied to every GRB, assigning a probability of a KNe association. This should enable dedicated deep searches for KNe counterparts associated with these GRBs, even without a GW observation. This will require developing a statistical framework (and a related analysis pipeline) that will carry out this exercise in a self-consistent manner. Moreover, these predictions will save the telescope time as it is impossible to perform a targeted search of all the GRBs, given the faint and rapid decay nature of KNe. The recent discovery of KNe associated with long GRBs further complicates the search, as previously thought to be exclusive to short GRBs.
@@ acknowledgments
We thank the Fermi/GBM team for providing extensive data through public archives. Dimple thanks P. Sreekumar, Ranjeev Mishra, Siddharth Pritam, Ben Gompertz for fruitful discussions. K.G.A. and Dimple acknowledge the Swarnajayanti grant DST/SJF/PSA-01/2017-18 of the Department of Science and Technology and SERB, India, and support from Infosys Foundation. KM and K.G.A. acknowledge the support from BRICS grant DST/ICD/BRICS/Call-5/CoNMuTraMO/2023 (G) funded by the Department of Science and Technology (DST), India.

### v2

@@ Abstract
Classification of gamma-ray bursts (GRBs) has been a long-standing puzzle in high-energy astrophysics. Recent observations challenge the traditional short vs. long viewpoint, where long GRBs are thought to originate from the collapse of massive stars and short GRBs from compact binary mergers. Machine learning (ML) algorithms have been instrumental in addressing this problem, revealing five distinct GRB groups within the Swift/BAT light curve data, two of which are associated with kilonovae (KNe). In this work, we extend our analysis to the Fermi/GBM catalog and identify five clusters using unsupervised ML techniques, consistent with the Swift/BAT results. These five clusters are well separated in fluence-duration plane, hinting at a potential link between fluence, duration and complexities (or structures) in the light curves of GRBs. Further, we confirm two distinct classes of KN-associated GRBs. The presence of GRB 170817A in one of the two KNe-associated clusters lends evidence to the hypothesis that this class of GRBs could potentially be produced by binary neutron star (BNS) mergers. The second KN-associated GRB cluster could potentially originate from NS-BH mergers. Future multimessenger observations of compact binaries in gravitational waves (GWs) and electromagnetic waves can be paramount in understanding these clusters better.
@@ 1 Introduction
Recent advances in machine learning (ML) and artificial intelligence (AI) are fostering a more detailed understanding of GRBs (Chattopadhyay & Maitra, 2017; Acuner & Ryde, 2018; Jespersen et al., 2020a; Salmon et al., 2022; Steinhardt et al., 2023; Garcia-Cifuentes et al., 2023; Mehta & Iyyani, 2024; Negro et al., 2024; Zhang et al., 2024; Kumar et al., 2024; Zhu et al., 2024). When applied to the analysis of prompt emission light curves, these techniques have revealed hidden patterns within the GRB population, suggesting the existence of subclasses beyond the traditional short–long GRB classification (Jespersen et al., 2020b; Steinhardt et al., 2023; Garcia-Cifuentes et al., 2023; Dimple et al., 2023a).
@@ 2.1 Data Reduction Pipeline
After standardization, we could recover 3349 GRBs from the Fermi/GBM catalog. For the ML algorithm, we used the data in three energy bands, two energy bands of NaI (8–50 keVand 50–300 keV) and one energy band of BGO (300–1000 keV) as NaI is more sensitive in the lower energy band, and BGO is more sensitive in the high energy bands. We stored the standardized data as a matrix sized [3349 x 81798], where each row represents a GRB and includes its light curves across three energy channels binned at 16 ms intervals.
@@ 2.2 Clustering using unsupervised learning
We generated the connectivity and density maps to further analyze the resulting low-dimensional representation. Connectivity maps visualize the probability of a connection between two data points, essentially depicting the aforementioned network in a two-dimensional format. On the other hand, density maps, along with the 2-dimensional embeddings, offer additional insights into the data structure. These maps help to understand the distribution and concentration of data points in a low-dimensional space. Finally, we identified clusters within the two-dimensional embedding using the AutoGMM module (Athey et al., 2019), which applies a Gaussian Mixture Model (GMM) to automatically determine the optimal number of clusters present in the data. AutoGMM utilizes parameters such as the minimum and maximum number of clusters for cluster identification. For our analysis, we set the minimum number of clusters to 2 and the maximum to 7. This range was chosen based on density maps, which visually guide the expected number of clusters in embeddings and help prevent overfitting. We used the Bayesian Information Criterion (BIC) to determine the best classification, maintaining default settings for other parameters.
Figure 2: GRB locations on fluence–$T_{\rm 90}$ plane: GRBs are color coded based on distinct clusters identified by AutoGMM. Notably, these clusters exhibit clear separation in the fluence-duration plane, suggesting that light curve features have dependence on GRB fluence and duration.
@@ 3.1 Distinct groups within GRBs
The consistency of five distinct clusters across multiple datasets and methodologies suggests that the GRB population exhibits diversity beyond just two classes. To further investigate these clusters, we examined their prompt emission properties. We observed that these clusters exhibit minimal differences in their spectral properties, such as $E_{\rm p}$ and $\alpha$ (See Figure A.1 in Appendix A). However, a clear trend in duration and fluence can be observed for these clusters. The separation of clusters based on their duration is expected, but fluence also contributes to distinguishing the clusters. For instance, clusters 2 and 4 overlap in their duration values but differ in fluence values, suggesting that both fluence and duration have influenced the clustering. To further explore these clusters, we plotted the GRBs from the five clusters on the fluence–$T_{\rm 90}$ plane (Figure 2), and interestingly, they are well separated in this plane. This allows the identification of various classes: short-duration bursts with low fluence (cluster 0), intermediate-duration bursts with low fluence (clusters 1 and 2), long-duration bursts with low fluence (cluster 3), and intermediate to long-duration bursts with high fluence (cluster 4).
It is important to note that even though we removed the fluence information during data standardization, fluence still plays a role in clustering. This suggests that fluence information is conveyed through the light curves themselves, either through duration or structures within the light curves. This implies that the complexities of light curves may depend on brightness, which is also suggested by Stern et al. (1997). Though the correlation between fluence and duration is well known (Horváth et al., 2001), our results hint that there can be a correlation between light curve shapes and brightness, which may have a deeper connection with the physics of the central engines of GRBs. A detailed study of the implications of this correlation for the central engines of GRBs may be the subject of future work.
@@
Given that our clustering is based solely on the prompt emission light curves, these distinct classes may represent different progenitors, central engines, emission mechanisms, or a combination of these. To date, evidence from optical bumps (SNe/KNe) suggests that the collapse of massive stars (collapsars) or the merging of compact objects (NS/BH/WD) are two possible scenarios for the origin of GRBs. We locate all the candidates for KNe and SNe on the connectivity maps. SN-associated GRBs are spread across the map, as evident from Figure 3. Interestingly, some of these candidates are adjacent to KN-associated GRBs. SN-associated GRBs 221009A, 130427A, lie very close to KN-associated GRBs 211211A and 230307A. Further, GRB 200826A, a short GRB associated with an SN, lies in the cluster shared by KN-associated GRBs, GRBs 170817A and 160821B along with one SN-associated GRB 141006A. Li et al. (2023) also reported similarity in the temporal and spectral properties of some of the KN- and SN-associated GRBs. This indicates that despite differences in the progenitors, they might have similar central engines and/or emission mechanisms, resulting in similar light curves. It is possible that collapsars and compact binary mergers may result in the formation of a central engine, such as a hyperaccreting black hole or a magnetar, which could be responsible for launching a relativistic jet (Duncan & Thompson, 1992; Usov, 1992; Popham et al., 1999; Narayan et al., 2001; Metzger, 2019).
We also located the magnetar giant flares (MGFs), which mimic the short GRBs. We have only one MGF in our sample, 200415A (see Figure 3), which is very close to GRB 200826A.
@@ 3.2 Two distinct populations of KN-associated GRBs
The results are presented in Figure 3, which interestingly confirms the existence of two clusters of KNe-associated GRBs based on Swift/BAT. Unlike Swift/BAT, Fermi/GBM has observed GRB 170817A, known to be a BNS merger from GW observations (Abbott et al., 2017) with a confirmed detection of a KN. Therefore, we find evidence favouring the conjecture of Dimple et al. (2023a) that the low fluence, small duration cluster may correspond to BNS mergers. Similarly, GRBs 170817A and 160821B are closely located in the same cluster, implying they could possibly originate from the same kind of progenitor systems. Dimple et al. (2023a) argued that the second KNe-associated cluster could potentially be arising from NS-BH mergers. As GRB 230307A is not detected by Swift/BAT, the proximity of GRBs 211211A and 230307A in the Fermi/GBM map suggests the similarity in their light curves and the progenitors as also suggested by Peng et al. (2024). Both bursts exhibit three consistent emission episodes: a precursor, a main burst, and an extended emission phase (see figure 1 in Peng et al. 2024). The spectral lags for corresponding episodes in both GRBs are nearly zero, and the minimum variability time scales are comparable (Peng et al., 2024). They are also lying in a close vicinity in the low-dimensional embeddings obtained using deep learning on Fermi/GBM data (Negro et al., 2024). However, multimessenger observation of an NS-BH system is necessary to confirm the prediction that GRBs in this cluster are products of NS-BH mergers. For the typical NS masses, the BH must be of mass $\sim 3M_{\odot}$ for the NS-BH to produce a GRB. Recent observation of an NS-BH merger GW230529 (Abac et al., 2024) with a BH between $2.5-4.5M_{\odot}$ indicates there may be a population of this low mass BH in the universe. With all these, the first multimessenger NS-BH merger will revolutionize our understanding of compact objects and GRBs.
However, the presence of SNe-associated GRBs in this cluster complicates this paradigm. As SNe-associated GRBs are likely to be associated with massive stellar explosions, why their light curves share features similar to those of the KNe-associated GRBs is unclear. If, for instance, the mass of the central engine plays an important role, then SN explosions that leave behind BHs of mass $2.5-4.5M_{\odot}$ may share similar features in the light curves as that of NSBH mergers whose BH mass also lies in the same range. Admittedly, this is only a speculation at present that needs a careful and dedicated study in the future.
@@
    Of these five clusters, two distinct clusters contain KNe-associated GRBs. One cluster contains GRB 170817A, which is associated with a BNS merger. This strengthens the case for these two clusters to be associated with BNS and NS-BH mergers as proposed by Dimple et al. (2023a). However, these clusters are surrounded by some SNe-associated GRBs, which may possibly indicate the universality of the central engines of GRBs regardless of their progenitors.
@@
    Despite the use of fluence-normalized light curves, the ML-based clusters are well-separated in the fluence-$T_{\rm 90}$ plane, hinting that fluence also influenced the clustering. This indicates that the light curve shapes and duration may have dependencies on the brightness (also suggested by Horváth et al. (2001) and Stern et al. (1997) earlier), and is worth exploring.
We next discuss the implications of these findings. Even though the Fermi and Swift datasets contain mostly distinct events, with only a small overlap, their clustering patterns are remarkably similar (see Appendix B and Figure B.1). This consistency highlights the robustness of the ML-based classification. Further, KNe-associated GRBs are tightly clustered at two locations in both datasets. These clustering patterns can be used for early identification of such GRBs. This may be achieved by reanalyzing the existing data with the light curve of a new GRB in question and asking whether it lies in one of the two clusters. Such an algorithm can be developed and applied to every GRB, assigning a probability of a KNe association. This should enable dedicated deep searches for KNe counterparts associated with these GRBs, even without a GW observation. This will require developing a statistical framework (and a related analysis pipeline) that will carry out this exercise in a self-consistent manner. Moreover, these predictions will save the telescope time as it is impossible to perform a targeted search of all the GRBs, given the faint and rapid decay nature of KNe. The recent discovery of KNe associated with long GRBs further complicates the search, as previously thought to be exclusive to short GRBs. However, some of these bursts are surrounded by SN-associated GRBs. It remains interesting to follow up on them, as detecting either a SN or a KN is valuable.
@@ acknowledgments
We extend our sincere thanks to the referee for the valuable comments, which enhanced the quality and clarity of the manuscript. We thank the Fermi/GBM and Swift/BAT team for providing extensive data through public archives. Dimple thanks P. Sreekumar, Ranjeev Mishra, Siddharth Pritam, Ben Gompertz for fruitful discussions. K.G.A. and Dimple acknowledge the Swarnajayanti grant DST/SJF/PSA-01/2017-18 of the Department of Science and Technology and SERB, India, and support from Infosys Foundation. KM and K.G.A. acknowledge the support from BRICS grant DST/ICD/BRICS/Call-5/CoNMuTraMO/2023 (G) funded by the Department of Science and Technology (DST), India. Dimple acknowledges MCNS for their warm hospitality during her stay in the campus.
@@ Appendix A Characteristics of the clusters
Appendix B Cross-Matching and Clustering Consistency of GRB Embeddings Across Swift and Fermi Datasets
We compared the embeddings obtained using the Fermi/GBM dataset with those obtained using the Swift/BAT dataset. To do this, we first updated our Swift/BAT sample (defined in Dimple et al. (2023a) with the GRB light curves observed up to September 2023. We used PCA-UMAP and AutoGMM to cluster the bursts in the Swift/BAT dataset. For this, we set the value for n_neighbors at 15 and min_dist at 0.01. We identified the clusters using AutoGMM, which identified 5 clusters in the dataset as shown in the right panel of Figure B.1. We cross-matched the clusters for common GRBs in the Swift and Fermi catalogs. We compared the cluster numbers (as identified in Figure B.1) of 375 common bursts and found the cluster number matched for 139 bursts. It is noteworthy that even those bursts which did not share the same cluster number in both the maps were seen to lie in the neighbouring clusters in most cases we visually inspected. For instance, Swift bursts in Cluster 0 (green) remained in the same cluster in the Fermi map, as seen in the figure. A similar trend can be seen in the case of Cluster 4 (red), which corresponds to the same cluster in the Fermi map, except for a few bursts, which are seen to spread slightly into the nearby clusters. The extent of mixing for other clusters is greater, where some of the bursts in the Fermi map are seen to move to nearby clusters w.r.t. the Swift map.
Potential reasons for these discrepancies may include several factors. One key factor could be the different energy ranges of the light curves used for clustering for two datasets. The two satellites overlap in the lower energy ranges, but Swift does not include the energy bands from 350-1000 keV, which are covered by Fermi. These higher energy bands may contain critical information, ultimately causing differences in the embeddings. Additionally, we used 16ms binning for Fermi/GBM light curves, which include more detailed structures. In contrast, the Swift/BAT light curves were in 64ms (as we could not recover many light curves with 16ms binning from the Swift/BAT catalog after data standardization). Therefore, we had more refined structures in the Fermi light curves compared to Swift, which can also be a factor for the differences in the embeddings for the two catalogs. The clustering analysis involved approximately 1500 GRBs in the Swift sample and around 3300 GRBs in the Fermi sample, with a small overlap of common GRBs. Given that clustering depends on the overall behavior of each sample, differences between the two samples are expected. Further, the analysis was conducted in the detector count space which might be influenced by the non-linear responses of the instruments. Furthermore, differences in the processing pipelines could lead to classifications being affected by data artifacts and decisions made during processing.
{mdframed}
Figure B.1: The figure displays the mapping of common GRBs, color-coded by their respective clusters, on the two-dimensional embeddings derived from the Swift and Fermi datasets.